# L(M)V-IQL: Multiple Intention Inverse Reinforcement Learning for Animal Behavior Characterization

### v1

@@ Abstract
In advancing the understanding of decision-making processes, mathematical models, particularly Inverse Reinforcement Learning (IRL), have proven instrumental in reconstructing animal’s multiple intentions amidst complex behaviors. Given the recent development of a continuous-time multi-intention IRL framework, there has been persistent inquiry into inferring discrete time-varying reward functions with multiple intention IRL approaches. To tackle the challenge, we introduce the Latent (Markov) Variable Inverse Q-learning (L(M)V-IQL) algorithms, a novel IRL framework tailored for accommodating discrete intrinsic rewards. Leveraging an Expectation-Maximization approach, we cluster observed trajectories into distinct intentions and independently solve the IRL problem for each. Demonstrating the efficacy of L(M)V-IQL through simulated experiments and its application to different real mouse behavior datasets, our approach surpasses current benchmarks in animal behavior prediction, producing interpretable reward functions. This advancement holds promise for neuroscience and psychology, contributing to a deeper understanding of animal decision-making and uncovering underlying brain mechanisms.
@@ 1 Introduction
The characterization of animal’s decision-making behavior stands as a fundamental objective within the field of behavioral neuroscience (Niv, 2009; Wilson & Collins, 2019). Prior research has formulated a variety of mathematical behavioral models across diverse tasks, encompassing generalized linear models and models based on reinforcement learning (Ashwood et al., 2022b; Beron et al., 2022). Such forward models facilitate the comprehension and comparison of decision-making strategies employed by both human and animal subjects. Additionally, they offer a low-dimensional behavioral representation suitable for regression analysis with neural activities (Hattori et al., 2019; Hamaguchi et al., 2022). Forward models require a empirically defined reward function that guides subjects optimizing their behavior during decision-making. However, defining a comprehensive and suitable reward function can pose challenges in complex behavioral tasks. Alyahyay et al. (2023) introduced a response-preparation task where subjects ought to hold a lever until a cue indicating the release signal. In this task, subjects can receive a binary extrinsic reward from the environment, whereas the intrinsic reward driving behavior such as hunger, thirst, engagement, associated with each timestamp is, however, not obvious to the experimenter. As another example, within a 127-node-labyrinth with a water port at the terminal, Rosenberg et al. (2021) observed that the navigation behavior of water-restricted mice is influenced not solely by the extrinsic water reward but also by intrinsic motivators, including their curiosity to explore the environment.
Inverse reinforcement learning (IRL) (Ng et al., 2000; Arora & Doshi, 2021) is a popular approach to recover a reward function that induces the observed behavior, assuming that the demonstrator was (softly) maximizing its long-term return. Along with the significant successes of IRL in autonomous driving (Kalweit et al., 2020; Nasernejad et al., 2023), robotics (Kumar et al., 2023; Chen et al., 2023), and healthcare domains (Coronato et al., 2020; Chan & van der Schaar, 2021), it appears to be emerging as an valuable tool for constructing mathematical behavior models in neuroscience research, as exemplified by Yamaguchi et al. (2018), Kwon et al. (2020), and Alyahyay et al. (2023). Classic IRL methods seek to identify a single, fixed reward function for a specific scenario. In contrast, Ashwood et al. (2022a) suggested that animal’s goals can evolve over time due to factors like fatigue, satiation, and curiosity. Under this assumption, they proposed the Dynamic Inverse Reinforcement Learning (DIRL) framework, which parametrizes the animal’s reward function as a time-varying linear combination of a small number of spatial reward maps, which are referred to as “goal maps”. By positing the existence of multiple goal maps with time-varying weights, DIRL allows the instantaneous reward function to vary continuously in time. This innovative framework achieved state-of-the-art performance in characterizing animal behavior. Nevertheless, persistent demands have emerged regarding an IRL framework incorporating discrete time-varying reward functions, particularly following the proposal by Ashwood et al. (2022b) that natural behaviors can be represented through a Markov chain characterized by alternating between discrete intentions.
To address this requirement, we propose the novel class of Latent (Markov) Variable Inverse Q-learning (L(M)V-IQL) algorithms, which extend the single reward Inverse Q-learning (IQL) framework from Kalweit et al. (2020) to solve IRL problems accounting for multiple intentions. We formulate an Expectation-Maximization (EM) approach to first cluster animal trajectories into multiple intentions, and then solving the IRL problem independently for each intention, respectively. We theoretically demonstrate that L(M)V-IQL can cover the most common two types of intention transition dynamics: generalized Bernoulli process and Markov process. Finally, we present the application of our framework in 1) a simulated Gridworld environment, 2) real mice navigation trajectories with known environment model from Rosenberg et al. (2021), and 3) real mice decision-making data from a dynamic two-armed bandit task with unknown environment model, showing that our methods outperform state-of-the-art in behavior prediction, demonstrate exceptional aptitude in capturing the intentions of animals, and provide interpretable reward functions corresponding to each identified intention.
2 Related Work
@@
Various approaches have been introduced to address multiple intention inverse reinforcement learning problems (Table 1). Notably, several frameworks based on parametric (Babes et al., 2011; Likmeta et al., 2021), or Bayesian non-parametric (Choi & Kim, 2012; Bighashdel et al., 2021) approaches share a similar objective to ours, allowing for multiple agents with distinct reward functions. However, these frameworks do not accommodate single agents with time-varying rewards. Gleave & Habryka (2018) developed a meta adversarial learning method for multi-task IRL problems. While their framework demonstrated high-level performance in real-world applications, it sacrifices theoretical interpretability and heavily relies on exploiting similarities between reward functions across tasks. Built on the Bayesian IRL approach, algorithms from Dimitrakakis & Rothkopf (2012), Michini & How (2012), and Surana & Srivastava (2014) posits that an agent’s trajectory can be divided into discrete behavioral states with corresponding unique reward function. Nevertheless all these algorithms assume linear reward functions and the Bayesian inference problem is intractable even for moderately sized finite-state IRL problems. Finally, the framework from Ashwood et al. (2022a) parametrizes the animal’s reward function as a time-varying linear combination of a small number of spatial reward maps with Gaussian random walk prior over weights, capturing continuous time-varying reward functions. Their approach pursue a related aim to ours, yet is limited to capturing continuous intra-episode variation of reward functions (as illustrated in Section 5.1), and difficult to adopt to other environments. Last but not least, most of the aforementioned algorithms are model-based, relying on a known transition dynamics of the environment, whereas in many scenarios, the environment model is unknown.
@@ 3.1 Inverse Reinforcement Learning
Consider a Markov Decision Process (MDP): $\{{\mathcal{S}},{\mathcal{A}},T,{\mathcal{R}},\gamma\}$, where ${\mathcal{S}}$ and ${\mathcal{A}}$ denotes the state- and action-space; $T\colon{\mathcal{S}}\times{\mathcal{A}}\to\Delta({\mathcal{S}})$ is the state transition function ($\Delta$ denotes the probability simplex) with $T(s,a,s^{\prime})\coloneqq\Pr(s^{\prime}\mid s,a)$; ${\mathcal{R}}\colon{\mathcal{S}}\times{\mathcal{A}}\to{\mathbb{R}}$ defines the reward function, and $\gamma\in[0,1)$ denotes the discount factor. The problem of inverse reinforcement learning is formally defined as:
Problem 1 (IRL problem).
Given the demonstration space ${\mathcal{D}}\coloneqq{\{\xi_{i}\}}_{i=1}^{N}$ with $N$ trajectories, where each trajectory is a sequence of state-action pairs, $\xi_{i}\coloneqq\{(s_{1},a_{1}),(s_{2},a_{2}),\ldots\}$, the IRL problem consists of finding a reward function ${\mathcal{R}}$ that maximizes the LL between agent demonstrations and the (soft) optimal policy $\pi_{{\mathcal{R}}}$ under ${\mathcal{R}}$:
$\operatorname*{maximize}_{{\mathcal{R}}}\quad\sum_{i=1}^{N}\log(\Pr(\xi_{i}\mid\pi_{{\mathcal{R}}}))\text{\,.}$ (1)
@@ 3.2 Inverse Q-learning
The class of Inverse Q-learning algorithms (Kalweit et al., 2020) provides a precise yet notably time-efficient solution to Problem 1, compared to the popular Maximum Entropy IRL algorithm from Ziebart et al. (2008) and some of its variants. It assumes that the demonstrations are collected from an agent following a Boltzmann policy according to its unknown optimal value function $Q^{*}$:
$\pi^{E}(s,a)\coloneqq\frac{\exp(Q^{*}(s,a))}{\sum_{A\in{\mathcal{A}}}\exp(Q^{*}(s,A))}\Rightarrow Q^{*}(s,a)=Q^{*}(s,b)+\log(\pi^{E}(s,a))-\log(\pi^{E}(s,b))\text{\,,}$ (2)
for all actions $a\in{\mathcal{A}}$ and $b\in{\mathcal{A}}_{\bar{a}}$ where ${\mathcal{A}}_{\bar{a}}\coloneqq{\mathcal{A}}\backslash\{a\}$. Using the Bellman optimality equation in Equation 2, the immediate reward of action $a$ in state $s$ can be expressed by the immediate reward of some other action $b\in{\mathcal{A}}_{\bar{a}}$, the respective log-probabilities and future action-values:
@@
where $r(s,a)\in{\mathcal{R}}$ is the unknown reward function, $d_{{\mathcal{A}}}$ denotes the dimension of ${\mathcal{A}}$, and $\eta_{s}^{a}\coloneqq\log(\pi^{E}(s,a))-\gamma\sum_{s^{\prime}\in{\mathcal{S}}}T(s,a,s^{\prime})\max_{a^{\prime}\in{\mathcal{A}}}Q^{*}(s^{\prime},a^{\prime})$. The resulting system of linear equations can be solved with least squares, leading to the model-based Inverse Action-value Iteration (IAVI) algorithm, which solves the IRL problem analytically in closed-form. To relax the assumption of an existing transition model and action probabilities, IAVI was further extended to the sampling-based model-free Inverse Q-learning (IQL) algorithm (Kalweit et al., 2020). They showed that the Boltzmann distribution induced by the optimal action-value function on the learned reward from IAVI and IQL is equivalent to the arbitrary demonstrated behavior distribution.
@@ 4 Inverse Q-learning about Multiple Intentions
We first define the Multiple Intention Inverse Reinforcement Learning (MI-IRL) problem accordingly:
Problem 2 (MI-IRL problem).
Let ${\mathcal{Z}}\coloneqq{\{z_{k}\}}_{k=1}^{K}$ be a $K$-dimensional latent state space with each $z_{k}\in{\mathcal{Z}}$ corresponding to one intention, and let ${\mathcal{D}}\coloneqq{\{\xi_{i}\}}_{i=1}^{N}$ be $N$ trajectories demonstrated by an agent each under one of the latent states without labels. The MI-IRL problem consists of inferring the latent state labels and the corresponding reward functions ${\{{\mathcal{R}}_{k}\}}_{k=1}^{K}$ in ${\mathcal{D}}$ such that under the $k$-th latent state the agent (softly) optimizes ${\mathcal{R}}_{k}$.
We adopt the EM (Dempster et al., 1977) as a straightforward approach to attack Problem 2. Let $\Theta$ be the set of parameters to be inferred, and let ${\mathcal{Y}}\coloneqq{\{y_{i}\}}_{i=1}^{N}$ be the set of latent state labels for each trajectory $\xi\in{\mathcal{D}}$, where $y_{i}=k$ if trajectory $i$ came from under latent state $z_{k}$. At iteration $\tau$ of the EM process before $\Theta$ converges, the expected value of the likelihood function ${\mathcal{L}}$ of $\Theta$ will be maximized as in the following update equation:
$\Theta^{\tau+1}\coloneqq\operatorname*{arg\,max}_{\Theta}\sum_{{\mathcal{Y}}}{\mathcal{L}}(\Theta\mid{\mathcal{D}},{\mathcal{Y}})\Pr({\mathcal{Y}}\mid{\mathcal{D}},\Theta^{\tau})$ (4)

### v2

# Multi-intention Inverse Q-learning for Interpretable Behavior Representation

@@ Abstract
In advancing the understanding of decision-making processes, Inverse Reinforcement Learning (IRL) have proven instrumental in reconstructing animal’s multiple intentions amidst complex behaviors. Given the recent development of a continuous-time multi-intention IRL framework, there has been persistent inquiry into inferring discrete time-varying rewards with IRL. To tackle the challenge, we introduce Latent (Markov) Variable Inverse Q-learning (L(M)V-IQL), a novel class of IRL algorthms tailored for accommodating discrete intrinsic reward functions. Leveraging an Expectation-Maximization approach, we cluster observed expert trajectories into distinct intentions and independently solve the IRL problem for each. Demonstrating the efficacy of L(M)V-IQL through simulated experiments and its application to different real mouse behavior datasets, our approach surpasses current benchmarks in animal behavior prediction, producing interpretable reward functions. This advancement holds promise for neuroscience and cognitive science, contributing to a deeper understanding of decision-making and uncovering underlying brain mechanisms.
Inverse Reinforcement Learning, Neuroscience, Decision-making
@@ 1 Introduction
G-EM-MLIRL (Nguyen et al., 2015) $\times$ linear known discrete
@@
Characterizing decision-making behavior stands as a fundamental objective within the field of behavioral neuroscience (Niv, 2009; Wilson & Collins, 2019). Prior research has formulated a variety of mathematical behavioral models across diverse tasks (Ashwood et al., 2022b; Beron et al., 2022), encompassing generalized linear models and models based on reinforcement learning. These forward models facilitate the comprehension and comparison of decision-making strategies employed by both human and animal subjects. Additionally, they offer a low-dimensional behavioral representation suitable for regression analysis with neural activities (Hattori et al., 2019; Hamaguchi et al., 2022). Forward models require an empirically defined reward function that guides subjects optimizing their behavior during decision-making. However, defining a comprehensive and suitable reward function can pose challenges in complex behavioral tasks. Alyahyay et al. (2023) introduced a response-preparation task where subjects ought to hold a lever until a cue indicating the release signal. In this task, subjects can receive a binary extrinsic reward from the environment, whereas the intrinsic reward driving behavior such as hunger, thirst, engagement, associated with each timestamp is, however, not obvious to the experimenter. As another example, within a 127-node-labyrinth with a water port at the terminal, Rosenberg et al. (2021) observed that the navigation behavior of water-restricted mice is influenced not solely by the extrinsic water reward but also by intrinsic motivators, including their curiosity to explore the environment.
Inverse reinforcement learning (IRL) (Ng et al., 2000; Arora & Doshi, 2021) is a popular approach to recover a reward function that induces the observed behavior, assuming that the demonstrator was maximizing its long-term return. Along with the significant successes of IRL in autonomous driving (Kalweit et al., 2020; Nasernejad et al., 2023), robotics (Kumar et al., 2023; Chen et al., 2023), and healthcare domains (Coronato et al., 2020; Chan & van der Schaar, 2021), it appears to be emerging as an valuable tool for constructing mathematical behavior models in neuroscience research, as exemplified by Yamaguchi et al. (2018), Kwon et al. (2020), and Alyahyay et al. (2023). Classic IRL methods seek to identify a single, fixed reward function for a specific scenario. In contrast, Ashwood et al. (2022a) suggested that animal’s goals can evolve over time due to factors like fatigue, satiation, and curiosity. Under this assumption, they proposed the Dynamic Inverse Reinforcement Learning (DIRL) framework, which parametrizes the animal’s reward function as a smoothed time-varying linear combination of a small number of spatial reward maps, which are referred to as ‘goal maps’. By positing the existence of multiple goal maps with time-varying weights, DIRL allows the instantaneous reward function to vary continuously in time. This innovative framework achieved state-of-the-art performance in animal behavior prediction. Nevertheless, persistent demands have emerged regarding an IRL framework incorporating discrete time-varying reward functions, particularly following the proposal by Ashwood et al. (2022b) that natural behaviors can be represented through a Markov chain characterized by alternating between discrete intentions.
To address this requirement, we propose the novel class of Latent (Markov) Variable Inverse Q-learning (L(M)V-IQL) algorithms, which extend the fixed-reward Inverse Q-learning (IQL) framework from Kalweit et al. (2020) to solve IRL problems accounting for multiple intentions. We formulate an Expectation-Maximization (EM) approach to first cluster animal trajectories into multiple intentions, and then solving the IRL problem independently for each intention. We theoretically demonstrate that L(M)V-IQL can cover the most common two types of intention transition dynamics: generalized Bernoulli process and Markov process. Finally, we present experiments on the application of our framework in 1) a simulated Gridworld environment; 2) real mice navigation trajectories with known environment model from the 127-node-labyrinth task (Rosenberg et al., 2021), serving as a benchmark for comparing with the state-of-the-art algorithm DIRL  (Ashwood et al., 2022a); and 3) real mice decision-making data with unknown environment model from a dynamic two-armed bandit task (De La Crompe et al., 2023). Demonstrating superior performance in behavior prediction, our methods showcase exceptional proficiency in capturing animals’ intentions through interpretable reward functions derived solely from their trajectories, surpassing the state-of-the-art approaches.
2 Related Work
@@
Various approaches have been introduced to address Multiple Intention Inverse Reinforcement Learning (MI-IRL) problems (Table 1). Notably, several frameworks based on parametric (Babes et al., 2011; Likmeta et al., 2021), or non-parametric (Choi & Kim, 2012; Bighashdel et al., 2021) approaches allow for learning from multiple agents with distinct reward functions. However, these frameworks do not accommodate single agents with time-varying rewards. Gleave & Habryka (2018) developed a meta adversarial learning method for multi-task IRL problems. While their framework demonstrated high-level performance in real-world applications, it sacrifices interpretability and heavily relies on exploiting similarities between reward functions across tasks.
In contrast, several approaches formulate MI-IRL problem as finding the maximum-likelihood partition of each trajectory, where each segment was generated from different locally consistent reward functions. To solve the problem, Dimitrakakis & Rothkopf (2012), Michini & How (2012), and Surana & Srivastava (2014) established Bayesian IRL frameworks, while Nguyen et al. (2015) proposed a probabilistic graphical model, generalizing the algorithm from Babes et al. (2011). The latter approach avoids the computationally intensive Bayesian inference problem, which is intractable even for moderately sized finite-state IRL problems. Nevertheless all these algorithms are confined to linearly parameterized reward functions, whereas our L(M)V-IQL can effectively learn non-linear rewards.
The state-of-the-art framework, DIRL (Ashwood et al., 2022a), parametrizes the expert’s reward function as a time-varying linear combination of a small number of non-linear spatial reward maps with Gaussian random walk prior over weights, capturing continuous time-varying rewards. Their approach pursues a related aim to ours, yet has the following limitations: 1) DIRL can only capture intra-episode variation of reward functions, while L(M)V-IQL can learn both intra- and inter-episode varying rewards, as discussed in Section 5.1. 2) According to Ashwood et al. (2022b), humans and animals may switch between multiple discrete strategies during perceptual decision-making. Such behavioral characteristics would be challenging for DIRL to capture, as it assumes a smooth Gaussian prior on the time-varying rewards. Since the objective function and problem solver of DIRL heavily rely on this prior (Ashwood et al., 2022a), switching to other less smoothed prior would not be trivial. This limitation strongly affects DIRL’s performance, as demonstrated in Section 5.2, where we compare L(M)V-IQL and DIRL on the same benchmark. 3) While DIRL excels at predicting mice navigation behavior in a labyrinth, it faces challenges when adapting to diverse behavioral tasks.
As with other parametric frameworks, L(M)V-IQL adopts a design choice that requires a prior input specifying the number of intentions. This deliberate decision enables precise customization and enhances the algorithm’s efficiency in scenarios where intention specification is clear and well-defined. Last but not the least, most of the aforementioned algorithms are model-based, relying on a known transition dynamics of the environment, whereas in many scenarios, the environment model is unknown. As an improvement, our algorithms can perform model-free learning, enabling their application in a wider range of environments.
@@ 3.1 Inverse Reinforcement Learning
Consider the following Markov Decision Process (MDP): $\{\mathcal{S},\mathcal{A},T,r,\gamma\}$, where $\mathcal{S}$ and $\mathcal{A}$ denotes the state-space and action-space, respectively; $T\colon\mathcal{S}\times\mathcal{A}\to\mathcal{P}(\mathcal{S})$ is the state transition function ($\mathcal{P}$ is the probability simplex) with $T(s,a,s^{\prime})\coloneqq\Pr(s^{\prime}\mid s,a)$; $r\colon\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ defines the reward function, and $\gamma\in[0,1)$ denotes the discount factor. Additionally, $\pi(s,a)\coloneqq\Pr(a\mid s)$ is used to represent the policy according to which actions are selected in the MDP. A formal definition of inverse reinforcement learning problems is then:
Problem 3.1 (IRL problem).
Given the demonstration space $\mathcal{D}\coloneqq{\{\xi_{i}\}}_{i=1}^{N}$ including $N$ trajectories provided by an expert in an MDP, where each $\xi_{i}\coloneqq\{(s_{1},a_{1}),(s_{2},a_{2}),\ldots\}$ is a sequence of state-action pairs, the IRL problem consists of finding a reward function $r$ that maximizes the log-likelihood between expert demonstrations and the optimal policy $\pi_{r}$ under $r$:
maximize (over $r$) $\sum_{i=1}^{N}\log(\Pr(\xi_{i}\mid\pi_{r}))\text{\,.}$ (1)
@@ 3.2 Inverse Q-learning
The class of Inverse Q-learning algorithms (Kalweit et al., 2020) provides a precise yet notably time-efficient solution to 3.1, compared to the popular Maximum Entropy IRL algorithm from Ziebart et al. (2008) and some of its variants. It assumes that the demonstrations are collected from an agent following a Boltzmann policy according to its unknown optimal value function $Q^{*}$:
$\pi^{E}(s,a)\coloneqq\frac{\exp(Q^{*}(s,a))}{\sum_{A\in\mathcal{A}}\exp(Q^{*}(s,A))}\text{\,.}$ (2)
Rearranging Equation 2 leads to $Q^{*}(s,a)=Q^{*}(s,b)+\log(\pi^{E}(s,a))-\log(\pi^{E}(s,b))$, for all actions $a\in\mathcal{A}$ and $b\in\mathcal{A}_{\bar{a}}$ where $\mathcal{A}_{\bar{a}}\coloneqq\mathcal{A}\backslash\{a\}$. Substituting the Bellman optimality equation, i.e. $Q^{*}(s,a)=r(s,a)+\gamma\sum_{s^{\prime}\in\mathcal{S}}T(s,a,s^{\prime})\max_{a^{\prime}\in\mathcal{A}}Q^{*}(s^{\prime},a^{\prime})$ (Howard, 1960), the immediate reward of action $a$ at state $s$ can be expressed by the immediate reward of some other action $b\in\mathcal{A}_{\bar{a}}$, the respective log-probabilities and future action-values:
@@
where $r(s,\cdot)$ is the unknown reward function at state $s\in\mathcal{S}$, $d_{\mathcal{A}}$ denotes the dimension of $\mathcal{A}$, and
$\eta_{s}^{a}\coloneqq\log(\pi^{E}(s,a))-\gamma\sum_{s^{\prime}\in\mathcal{S}}T(s,a,s^{\prime})\max_{a^{\prime}\in\mathcal{A}}Q^{*}(s^{\prime},a^{\prime})\text{\,.}$ (4)
The resulting system of linear equations can be solved with least squares, leading to the model-based Inverse Action-value Iteration (IAVI) algorithm, which solves the IRL problem analytically in closed-form. To relax the assumption of an existing transition model and action probabilities, IAVI was further extended to the sampling-based model-free Inverse Q-learning (IQL) algorithm (Kalweit et al., 2020). They showed that the Boltzmann distribution induced by the optimal action-value function on the learned reward from IAVI and IQL is equivalent to the arbitrary demonstrated behavior distribution.
@@ 4 Inverse Q-learning about Multiple Intentions
We first provide a formal definition of MI-IRL problems accordingly:
Problem 4.1 (MI-IRL problem).
Let $\mathcal{Z}\coloneqq{\{z_{k}\}}_{k=1}^{K}$ be a $K$-dimensional latent state space with each $z_{k}\in\mathcal{Z}$ corresponding to one intention, and let $\mathcal{D}\coloneqq{\{\xi_{i}\}}_{i=1}^{N}$ be $N$ trajectories demonstrated by an agent each under one of the latent states without labels. The MI-IRL problem consists of inferring the latent state labels and the corresponding reward functions ${\{r_{k}\}}_{k=1}^{K}$ in $\mathcal{D}$ such that under the $k$-th latent state the agent (softly) optimizes $r_{k}$.
We adopt the EM (Dempster et al., 1977) as a straightforward approach to attack 4.1. Let $\Theta$ be the set of parameters to be inferred, and let $\mathcal{Y}\coloneqq{\{y_{i}\}}_{i=1}^{N}$ be the set of latent state labels for each trajectory $\xi\in\mathcal{D}$, where $y_{i}=k$ if trajectory $i$ came from under latent state $z_{k}$. At iteration $\tau$ of the EM process before $\Theta$ converges, the expected value of the likelihood function $\mathcal{L}$ of $\Theta$ will be maximized as in the following update equation:
$\Theta^{\tau+1}\coloneqq\operatorname*{argmax}_{\Theta}\sum_{\mathcal{Y}}\mathcal{L}(\Theta\mid\mathcal{D},\mathcal{Y})\Pr(\mathcal{Y}\mid\mathcal{D},\Theta^{\tau})\text{\,.}$ (5)